The area of a rectangle is 64cm squared. What are the dimensions if the perimeter is a minimum?
Answer by Boreal(15235)

length is x
width is y
xy=64 cm^2
y=64/x
perimeter is 2x+2y=2x+128/x
That has to be a minimum
derivative of that is 2-128/x^2
Set that equal to 0
128/x^2=2
2x^2=128
x=8 cm
y=8 cm
It is a square, and the perimeter is a minimum of 32 cm.